Soy bean meals are 14% protien; cornmeal is 7% protien How many pounds of each should be mixed together to get a 280lb mixture that is 10% protien?
Answer by robertb(4012)

0.14x + 0.07(280 - x) = 0.10*280
==> 0.14x + 19.6 - 0.07x = 28
==> 0.07x = 8.4
==> x = 120 pounds of soy bean meal, and 280 - 120 = 160 pounds of corn meal.